# FLoC: Facility Location-Based Efficient Visual Token Compression for Long Video Understanding

### v1

@@ Abstract
Recent studies in long video understanding have harnessed the advanced visual-language reasoning capabilities of Large Multimodal Models (LMMs), driving the evolution of video-LMMs specialized for processing extended video sequences. However, the scalability of these models is severely limited by the overwhelming volume of visual tokens generated from extended video sequences. To address this challenge, this paper proposes FLoC, an efficient visual token compression framework based on the facility location function, a principled approach that swiftly selects a compact yet highly representative and diverse subset of visual tokens within a predefined budget on the number of visual tokens. By integrating the lazy greedy algorithm, our method achieves remarkable efficiency gains by swiftly selecting a compact subset of tokens, drastically reducing the number of visual tokens while guaranteeing near-optimal performance. Notably, our approach is training-free, model-agnostic, and query-agnostic, providing a versatile solution that seamlessly integrates with diverse video-LLMs and existing workflows. Extensive evaluations on large-scale benchmarks, such as Video-MME, MLVU, and LongVideoBench, demonstrate that our framework consistently surpasses recent compression techniques, highlighting not only its effectiveness and robustness in addressing the critical challenges of long video understanding, but also its efficiency in processing speed.
@@ 1 Introduction
With the recent emergence of Large Language Models (LLMs) in natural language processing, there has been a surge of interest in extending their capabilities to the visual domain Achiam et al. (2023). By utilizing the visual embeddings as token inputs to the LLMs, referred to as visual tokens, these Large Multimodal Models (LMMs) have already demonstrated their performances surpassing human-level accuracy on vision tasks, such as visual question answering Liu et al. (2024); Fang et al. (2024); Team et al. (2023). More recently, the research focus has shifted towards enabling these models to understand video sequences Lin et al. (2023), giving rise to video-LMMs Song et al. (2024); Xue et al. (2024); Wang et al. (2024a); Balazevic et al. (2024). Such models not only excel in tasks like captioning Krishna et al. (2017); Xu et al. (2015); Vinyals et al. (2015), event detection Xu et al. (2019); Shou et al. (2021), and action recognition Zhao et al. (2017); Simonyan & Zisserman (2014), but also show significant potential in various real-world applications, including surveillance through CCTV systems, immersive experiences in smart glasses, and autonomous navigation for mobile robots.
@@
Despite this progress, long video understanding remains particularly challenging due to the explosive growth in the number of visual tokens as the video sequence length increases Xue et al. (2024); Fu et al. (2024). When dealing with high-resolution or long-duration videos (e.g., 4K content), it becomes computationally infeasible to process every token end-to-end, especially given that most LLM-based architectures support input contexts of only 4K to 32K tokens. This limitation is exacerbated in real-world scenarios: for instance, continuous CCTV footage can span days or weeks, smart glasses may capture extended, first-person video streams, and mobile robots frequently operate in dynamic environments requiring real-time video analysis. Consequently, the gap between human-level performance and current model capabilities still exists, highlighting the complexity and significance of this research direction.
@@
To tackle the issue of handling long video sequences, visual token compression is indispensable. In practice, when examining consecutive frames of a video, many tokens share highly redundant information unless there is a substantial scene change Potapov et al. (2014). Eliminating these redundancies often does not harm the downstream performance, while excessively pruning tokens could lead to the loss of critical information. It is therefore critical to strike a delicate balance in token compression to minimize information loss.
@@
Previous approaches to selecting representative visual tokens often relied on filtering out temporally redundant tokens or frames Shen et al. (2024); Tao et al. (2024) or clustering techniques to extract representative information from each cluster Wang et al. (2024b); Shang et al. (2024); Zhang et al. (2024a). While these methods may work at a reasonable level, they often fall short in capturing the full diversity needed to interpret complex visual scenes. Consider a scenario where a user wearing smart glasses searches for car keys in a cluttered room. Visual tokens representing the small object of interest (the keys) occur infrequently and sparsely within the video sequence, whereas tokens depicting general scenery, such as furniture or background, appear repeatedly and redundantly. In this setting, clustering-based approaches are likely to fail in capturing rare but important tokens—such as those corresponding to the keys—since they primarily focus on densely populated regions in the feature space. Therefore, a visual token compression algorithm that simultaneously ensures representativeness and diversity is essential to effectively retain these critical but sparse visual cues.
@@
In order to overcome these limitations, we propose a novel visual token compression algorithm based on the facility location function Lin & Bilmes (2011); Lin et al. (2009). Our approach interprets token selection through the lens of submodular optimization, ensuring that the selected set of tokens covers all original tokens under a given budget constraint. Specifically, each subset considers the similarity between its subset and the entire tokens, enabling to include diverse information of the entire video sequence. While finding the optimal subsets in this manner is known to be a NP-hard problem, we sidestep the computational overhead by utilizing the lazy greedy algorithm Minoux (1978), enabling to select the visual tokens with minimal computational overheads. As a result, the chosen tokens are both representative and diverse, effectively preserving essential information for video understanding tasks. Our experiments on benchmarks such as Video-MME, and LongVideoBench Zhou et al. (2024); Wu et al. (2025) demonstrate the superiority of our method over existing approaches.
@@ 2 Related Work
Sampling / Pooling A common and straightforward strategy to deal with the abundance of visual tokens in long video sequences is to reduce the input size via pooling or sampling Potapov et al. (2014); Cai et al. (2024); Qu et al. (2024); Wu (2024). For instance, uniform sampling of frames or pooling across spatial/temporal dimensions can substantially cut down the computational overhead and memory usage. However, these methods often ignore the semantic importance of certain frames or regions. Such a one-size-fits-all approach may discard critical cues or overly compress redundant segments, leading to suboptimal performance when higher-level understanding of video content is required.
@@
Clustering Another widely studied line of research involves clustering techniques to group similar frames or tokens and select representative exemplars de Avila et al. (2011); Khosla et al. (2013); Wang et al. (2024b); Shang et al. (2024); Zhang et al. (2024a). By partitioning the visual space into clusters, these methods attempt to capture the overall distribution of the video content, retaining only the most “central” examples in each cluster. While clustering can better preserve representativeness than naive sampling, it can still struggle to guarantee coverage of rare but potentially important events. Moreover, the offline clustering process may be computationally expensive, especially for long videos, and is typically not optimized in an end-to-end manner, which can result in mismatches between clustering objectives and downstream video understanding tasks.
@@
Query-Aware Compression In query-aware or task-specific compression, the aim is to select those frames or tokens that are most relevant to a given query, user interest, or downstream task Zhang et al. (2016); Shen et al. (2024); Korbar et al. (2024); Wang et al. (2024b). This category of methods can effectively reduce the search space by focusing on what is deemed important. However, they require prior knowledge of the query or task, making them less flexible for general-purpose or zero-shot scenarios. When the query space expands or changes, such approaches often need retraining or redesign, limiting their applicability in dynamic environments (e.g., surveillance systems, smart glasses, or robots) where the set of possible queries is not fixed.
@@
Retraining Learnable compression algorithms employ neural networks to decide which tokens or frames to discard or keep Zhang et al. (2025); Argaw et al. (2024); Lee et al. (2025). By training end-to-end, they can theoretically capture complex patterns and adapt to different tasks. Nonetheless, these methods tend to require large labeled datasets and substantial training time. They are also dependent on model architecture and specific training objectives, which makes them less model-agnostic. Consequently, deploying such methods in rapidly evolving research fields or on resource-constrained platforms (e.g., embedded systems in mobile robots) can be challenging.
@@
In contrast to the above approaches, our method operates in a training-free, plug-and-play fashion, allowing it to be easily integrated into existing pipelines with minimal overhead. Built on the principle of facility location Lin & Bilmes (2011); Lin et al. (2009), it interprets token selection as a submodular optimization problem, ensuring both representativeness and diversity under a given budget constraint. Additionally, we adopt a lazy greedy algorithm that significantly reduces computation time while maintaining near-optimal performance Minoux (1978). By decoupling the compression strategy from the underlying vision model, our approach remains model-agnostic, thus enabling seamless deployment in various real-world scenarios, from large-scale video analytics to on-device processing for surveillance, smart eyewear, and mobile robots. Moreover, our proposed approach operates in a query-agnostic manner, independent of user input. Unlike query-aware methods that require recompression for each incoming query and must retain all uncompressed tokens in memory, our method performs a one-time compression and stores only the compressed tokens. This leads to significant gains in both computational and memory efficiency.
@@ 3.2 Submodular Facility Location Function
We utilize the facility location function Lin & Bilmes (2011); Lin et al. (2009), a widely adopted submodular function, to select a representative and diverse subset of visual tokens. Formally, given a ground set $V$ of visual tokens, the facility location objective is defined as follows:
@@
Finding an optimal subset that maximizes the facility location function is known to be NP-hard. To address this complexity, a common approximation method is the greedy algorithm, which iteratively selects tokens with the highest marginal gain until the budget constraint is satisfied. This greedy selection method guarantees a solution with a performance lower bound of $(1-1/e)\approx 0.632$ relative to the optimal solution Nemhauser et al. (1978). Specifically, the greedy algorithm incrementally adds the token that provides the largest increase in coverage at each iteration. To further enhance computational efficiency, we implement a lazy greedy algorithm Minoux (1978), which significantly reduces the computational overhead by postponing the update of marginal gains until absolutely necessary, thus allowing efficient subset selection even for very large token sets.
The lazy greedy algorithm significantly reduces computational complexity compared to the naive greedy approach. While the naive greedy algorithm for maximizing submodular functions has a time complexity of $O(nK)$, the lazy greedy approach leverages the submodularity property to avoid unnecessary recomputation of marginal gains. By using a priority queue, it updates marginal gains only when needed, achieving empirical speedups often approaching an order of magnitude. Consequently, it becomes particularly efficient for handling numerous visual tokens and enabling real-time processing of long videos. Algorithm 1 presents the detailed procedure of our proposed method, and Figure 3 provides a toy example illustrating how this method works.
@@
Comp.
Ratio
Tokens Method
Video
MME
MLVU LVB Avg. Tokens Method
Video
MME
MLVU LVB Avg.
1 24576 - 66.33 70.31 60.51 65.72 16384 - 66.63 72.68 59.39 66.23
$2^{-3}$ 3072 TS-LLaVA 61.15 67.57 55.20 61.31 2048 TS-LLaVA 62.78 67.30 56.02 62.03
LongVU 62.19 66.61 55.42 61.41 LongVU 64.70 69.50 55.35 63.18
DivPrune 61.63 67.57 56.17 61.79 DivPrune 64.07 70.06 56.92 63.68
Random 60.30 66.24 55.72 60.75 Random 60.59 65.69 56.02 60.77
DyCoke 62.11 67.53 55.12 61.59 DyCoke 63.96 68.45 55.72 62.71
PruneVid 58.19 64.54 54.15 58.96 PruneVid 57.41 62.05 53.48 57.65
@@
$2^{-4}$ 1536 TS-LLaVA 58.78 64.67 52.51 58.65 1024 TS-LLaVA 59.63 64.95 53.85 59.48
LongVU 58.07 62.97 52.73 57.92 LongVU 56.48 60.12 51.31 55.97
DivPrune 58.85 64.67 54.00 59.17 DivPrune 61.93 68.08 54.82 61.61
Random 57.44 63.80 53.63 58.29 Random 59.74 64.77 54.23 59.58
DyCoke 57.00 63.02 53.78 57.93 DyCoke 61.37 65.13 53.10 59.87
PruneVid 54.11 61.59 51.83 55.84 PruneVid 53.81 59.48 52.28 55.19
@@
$2^{-5}$ 768 TS-LLaVA 55.07 62.37 50.49 55.98 512 TS-LLaVA 58.89 63.89 53.33 58.70
LongVU 53.41 58.42 50.34 54.06 LongVU 55.96 59.52 51.01 55.50
DivPrune 55.78 61.91 52.28 56.66 DivPrune 60.85 65.46 52.88 59.73
Random 55.56 61.41 49.89 55.62 Random 57.30 63.57 52.13 57.67
DyCoke 54.37 59.98 51.38 55.24 DyCoke 59.22 62.60 51.98 57.93
PruneVid 51.11 58.51 49.66 53.09 PruneVid 51.41 56.39 49.96 52.59
@@
4 Experiments
4.1 Models
Qwen2.5-VL Bai et al. (2025) is an advanced vision-language model capable of handling high-resolution images and long video sequences. It introduces dynamic resolution processing via a Window Attention-based Vision Transformer and supports absolute temporal encoding.
InternVL3 Zhu et al. (2025) is a multimodal model designed with native vision-language pretraining and Cascade Reinforcement Learning. For long video understanding, it incorporates a Visual Resolution Router to dynamically allocate visual token capacity across frames.
Others. We also conducted experiments on Qwen2-VL Wang et al. (2024a) and LLaVA-Next-Video Zhang et al. (2024c) models to further validate the generalizability of our approach. Due to space limitations, detailed results and analysis for these models are provided in the Appendix.  [^1] Qwen2.5-VL, Qwen2-VL, and LLaVA-Video-7B-Qwen2 are all under the Apache-2.0 license. InternVL3 is under the MIT license.

### v2

@@ Abstract
Recent studies in long video understanding have harnessed the advanced visual-language reasoning capabilities of Large Multimodal Models (LMMs), driving the evolution of video-LMMs specialized for processing extended video sequences. However, the scalability of these models is severely limited by the overwhelming volume of visual tokens generated from extended video sequences. To address this challenge, we propose FLoC, an efficient visual token compression framework based on the facility location function, a principled approach that swiftly selects a compact yet highly representative and diverse subset of visual tokens within a predefined budget on the number of visual tokens. By integrating the lazy greedy algorithm, our method achieves remarkable efficiency gains by swiftly selecting a compact subset of tokens, drastically reducing the number of visual tokens while guaranteeing near-optimal performance. Notably, our approach is training-free, model-agnostic, and query-agnostic, providing a versatile solution that seamlessly integrates with diverse video-LLMs and existing workflows. Extensive evaluations on large-scale benchmarks, such as Video-MME, MLVU, LongVideoBench, and EgoSchema, show that our framework consistently surpasses recent compression techniques, highlighting its effectiveness and robustness in addressing the challenges of long video understanding as well as its processing efficiency.
@@ 1 Introduction
With the recent emergence of Large Language Models (LLMs) in natural language processing, there has been a surge of interest in extending their capabilities to the visual domain (Achiam et al., 2023). By utilizing the visual embeddings as token inputs to the LLMs, referred to as visual tokens, these Large Multimodal Models (LMMs) have already demonstrated their performances surpassing human-level accuracy on vision tasks, such as visual question answering (Liu et al., 2024; Fang et al., 2024; Team et al., 2023). More recently, the research focus has shifted towards enabling these models to understand video sequences (Lin et al., 2023), giving rise to video-LMMs (Song et al., 2024; Xue et al., 2024; Wang et al., 2024a; Balazevic et al., 2024). Such models not only excel in tasks like captioning (Krishna et al., 2017; Xu et al., 2015; Vinyals et al., 2015), event detection (Xu et al., 2019; Shou et al., 2021), and action recognition (Zhao et al., 2017; Simonyan and Zisserman, 2014), but also show significant potential in various real-world applications, including surveillance through CCTV systems, immersive experiences in smart glasses, and autonomous navigation for mobile robots.
@@
Despite this progress, long video understanding remains particularly challenging due to the explosive growth in the number of visual tokens as the video sequence length increases (Xue et al., 2024; Fu et al., 2024). When dealing with high-resolution or long-duration videos (e.g., 4K content), it becomes computationally infeasible to process every token end-to-end, especially given that most LLM-based architectures support input contexts of only 4K to 32K tokens. This limitation is exacerbated in real-world scenarios: for instance, continuous CCTV footage can span days or weeks, smart glasses may capture extended, first-person video streams, and mobile robots frequently operate in dynamic environments requiring real-time video analysis. Consequently, the gap between human-level performance and current model capabilities still exists, highlighting the complexity and significance of this research direction.
@@
To tackle the issue of handling long video sequences, visual token compression is indispensable. In practice, when examining consecutive frames of a video, many tokens share highly redundant information unless there is a substantial scene change (Potapov et al., 2014). Eliminating these redundancies often does not harm the downstream performance, while excessively pruning tokens could lead to the loss of critical information. It is therefore critical to strike a delicate balance in token compression to minimize information loss.
@@
Previous approaches to selecting representative visual tokens often relied on filtering out temporally redundant tokens or frames (Shen et al., 2024; Tao et al., 2024) or clustering techniques to extract representative information from each cluster (Wang et al., 2024b; Shang et al., 2024; Zhang et al., 2024a). While these methods may work at a reasonable level, they often fall short in capturing the full diversity needed to interpret complex visual scenes. Consider a scenario where a user wearing smart glasses searches for car keys in a cluttered room. Visual tokens representing the small object of interest (the keys) occur infrequently and sparsely within the video sequence, whereas tokens depicting general scenery, such as furniture or background, appear repeatedly and redundantly. In this setting, clustering-based approaches are likely to fail in capturing rare but important tokens—such as those corresponding to the keys—since they primarily focus on densely populated regions in the feature space. Therefore, a visual token compression algorithm that simultaneously ensures representativeness and diversity is essential to effectively retain these critical but sparse visual cues.
@@
In order to overcome these limitations, we propose a novel visual token compression algorithm based on the facility location function (Lin and Bilmes, 2011; Lin et al., 2009). Our approach interprets token selection through the lens of submodular optimization, ensuring that the selected set of tokens covers all original tokens under a given budget constraint. Specifically, each subset considers the similarity between its subset and the entire tokens, enabling to include diverse information of the entire video sequence. While finding the optimal subsets in this manner is known to be a NP-hard problem, we sidestep the computational overhead by utilizing the lazy greedy algorithm (Minoux, 1978), enabling to select the visual tokens with minimal computational overheads. As a result, the chosen tokens are both representative and diverse, effectively preserving essential information for video understanding tasks. Our experiments on benchmarks such as Video-MME, MLVU, LongVideoBench, and EgoSchema  (Fu et al., 2024; Mangalam et al., 2023; Zhou et al., 2024; Wu et al., 2025) demonstrate the superiority of our method over existing approaches.
@@ 2 Related Work
Sampling / Pooling A common and straightforward strategy to deal with the abundance of visual tokens in long video sequences is to reduce the input size via pooling or sampling (Potapov et al., 2014; Cai et al., 2024; Qu et al., 2024; Wu, 2024). For instance, uniform sampling of frames or pooling across spatial/temporal dimensions can substantially cut down the computational overhead and memory usage. However, these methods often ignore the semantic importance of certain frames or regions. Such a one-size-fits-all approach may discard critical cues or overly compress redundant segments, leading to suboptimal performance when higher-level understanding of video content is required.
@@
Clustering Another widely studied line of research involves clustering techniques to group similar frames or tokens and select representative exemplars (de Avila et al., 2011; Khosla et al., 2013; Wang et al., 2024b; Shang et al., 2024; Zhang et al., 2024a). By partitioning the visual space into clusters, these methods attempt to capture the overall distribution of the video content, retaining only the most “central” examples in each cluster. While clustering can better preserve representativeness than naive sampling, it can still struggle to guarantee coverage of rare but potentially important events. Moreover, the offline clustering process may be computationally expensive, especially for long videos, and is typically not optimized in an end-to-end manner, which can result in mismatches between clustering objectives and downstream video understanding tasks.
@@
Query-Aware Compression In query-aware or task-specific compression, the aim is to select those frames or tokens that are most relevant to a given query, user interest, or downstream task (Zhang et al., 2016; Shen et al., 2024; Korbar et al., 2024; Wang et al., 2024b). This category of methods can effectively reduce the search space by focusing on what is deemed important. However, they require prior knowledge of the query or task, making them less flexible for general-purpose or zero-shot scenarios. When the query space expands or changes, such approaches often need retraining or redesign, limiting their applicability in dynamic environments (e.g., surveillance systems, smart glasses, or robots) where the set of possible queries is not fixed.
@@
Retraining Learnable compression algorithms employ neural networks to decide which tokens or frames to discard or keep (Zhang et al., 2025; Argaw et al., 2024; Lee et al., 2025). By training end-to-end, they can theoretically capture complex patterns and adapt to different tasks. Nonetheless, these methods tend to require large labeled datasets and substantial training time. They are also dependent on model architecture and specific training objectives, which makes them less model-agnostic. Consequently, deploying such methods in rapidly evolving research fields or on resource-constrained platforms (e.g., embedded systems in mobile robots) can be challenging.
@@
In contrast to the above approaches, our method operates in a training-free, plug-and-play fashion, allowing it to be easily integrated into existing pipelines with minimal overhead. Built on the principle of facility location (Lin and Bilmes, 2011; Lin et al., 2009), it interprets token selection as a submodular optimization problem, ensuring both representativeness and diversity under a given budget constraint. Additionally, we adopt a lazy greedy algorithm that significantly reduces computation time while maintaining near-optimal performance (Minoux, 1978). By decoupling the compression strategy from the underlying vision model, our approach remains model-agnostic, thus enabling seamless deployment in various real-world scenarios, from large-scale video analytics to on-device processing for surveillance, smart eyewear, and mobile robots. Moreover, our proposed approach operates in a query-agnostic manner, independent of user input. Unlike query-aware methods that require recompression for each incoming query and must retain all uncompressed tokens in memory, our method performs a one-time compression and stores only the compressed tokens. This leads to significant gains in both computational and memory efficiency.
@@ 3.2 Submodular Facility Location Function
We utilize the facility location function (Lin and Bilmes, 2011; Lin et al., 2009), a widely adopted submodular function, to select a representative and diverse subset of visual tokens. Formally, given a ground set $V$ of visual tokens, the facility location objective is defined as follows:
@@
Finding an optimal subset that maximizes the facility location function is known to be NP-hard. To address this complexity, a common approximation method is the greedy algorithm, which iteratively selects tokens with the highest marginal gain until the budget constraint is satisfied. This greedy selection method guarantees a solution with a performance lower bound of $(1-1/e)\approx 0.632$ relative to the optimal solution (Nemhauser et al., 1978). Specifically, the greedy algorithm incrementally adds the token that provides the largest increase in coverage at each iteration.
To further enhance computational efficiency, we implement a lazy greedy algorithm (Minoux, 1978), which significantly reduces the computational overhead by avoiding unnecessary recomputation of marginal gains. Specifically, the algorithm exploits the submodularity (diminishing returns) property of the facility location function $f$. Formally, for any subsets $A\subseteq B\subseteq V$ and a token $v\in V\setminus B$, the marginal gain satisfies: f(A ∪{v}) - f(A) ≥f(B ∪{v}) - f(B) This inequality implies that the marginal benefit of adding a visual token $v$ can only decrease or remain constant as the selected subset grows. Consequently, the marginal gain computed in a previous iteration serves as a valid upper bound for the current marginal gain. We leverage this by maintaining a priority queue of these upper bounds. In each step of the search process, we pop the candidate $v^{*}$ with the highest upper bound and recompute its exact marginal gain $\delta$ with respect to the current subset. If $\delta$ remains greater than or equal to the upper bounds of all other candidates in the queue, submodularity guarantees that $v^{*}$ is the optimal choice for the current iteration without needing to re-evaluate the rest. Algorithm 1 outlines the detailed procedure, and Figure 3 provides a visual illustration of this process.
The lazy greedy algorithm significantly reduces computational complexity compared to the naive greedy approach. While the naive greedy algorithm for maximizing submodular functions has a time complexity of $O(nK)$, the lazy greedy approach leverages the submodularity property to avoid unnecessary recomputation of marginal gains. By using a priority queue, it updates marginal gains only when needed, achieving empirical speedups often approaching an order of magnitude. Consequently, it becomes particularly efficient for handling numerous visual tokens and enabling real-time processing of long videos.
@@
4 Experiments
4.1 Models
Qwen2.5-VL (Bai et al., 2025) is an advanced vision-language model capable of handling high-resolution images and long video sequences. It introduces dynamic resolution processing via a Window Attention-based Vision Transformer and supports absolute temporal encoding.
InternVL3 (Zhu et al., 2025) is a multimodal model designed with native vision-language pretraining and Cascade Reinforcement Learning. For long video understanding, it incorporates a Visual Resolution Router to dynamically allocate visual token capacity across frames.
Others. We also conducted experiments on Qwen2-VL (Wang et al., 2024a) and LLaVA-Next-Video (Zhang et al., 2024c) models to further validate the generalizability of our approach. Due to space limitations, detailed results and analysis for these models are provided in the Appendix.  [^1] Qwen2.5-VL, Qwen2-VL, and LLaVA-Video-7B-Qwen2 are all under the Apache-2.0 license. InternVL3 is under the MIT license.
@@ 4.1 Models
Comp. Ratio Method Video MME MLVU LVB Ego Schema Avg. Video MME MLVU LVB Ego Schema Avg.
1 - 66.33 70.31 60.51 61.40 64.64 66.63 72.68 59.39 70.00 67.18
$2^{-3}$ TS-LLaVA 61.15 67.57 55.20 59.60 60.88 62.78 67.30 56.02 68.20 63.58
LongVU 62.19 66.61 55.42 59.40 60.91 64.70 69.50 55.35 69.20 64.69
DivPrune 61.63 67.57 56.17 58.40 60.94 64.07 70.06 56.92 65.00 64.01
Random 60.30 66.24 55.72 58.60 60.22 60.59 65.69 56.02 65.20 61.88
DyCoke 62.11 67.53 55.12 59.60 61.09 63.96 68.45 55.72 69.00 64.28
PruneVid 58.19 64.54 54.15 54.20 57.77 57.41 62.05 53.48 62.80 58.94
STTM 59.52 63.57 54.60 55.80 58.37 63.52 64.26 54.90 66.20 62.22
Scissor 58.59 65.04 54.08 56.40 58.53 61.15 67.76 55.12 65.80 62.46
FastVID 60.89 67.31 57.14 58.60 60.99 - - - - -
@@
$2^{-4}$ TS-LLaVA 58.78 64.67 52.51 57.20 58.29 59.63 64.95 53.85 62.80 60.31
LongVU 58.07 62.97 52.73 55.40 57.29 56.48 60.12 51.31 60.40 57.08
DivPrune 58.85 64.67 54.00 55.80 58.33 61.93 68.08 54.82 61.80 61.66
Random 57.44 63.80 53.63 58.20 58.27 59.74 64.77 54.23 66.60 61.34
DyCoke 57.00 63.02 53.78 54.00 56.95 61.37 65.13 53.10 67.40 61.75
PruneVid 54.11 61.59 51.83 52.00 54.88 53.81 59.48 52.28 58.40 55.99
STTM 57.15 61.73 51.68 50.80 55.34 60.15 62.93 52.28 63.40 59.69
Scissor 55.26 60.95 53.55 54.00 55.94 58.89 64.44 53.77 63.40 60.13
FastVID 58.67 65.52 54.23 57.20 58.91 - - - - -
@@
$2^{-5}$ TS-LLaVA 55.07 62.37 50.49 54.60 55.63 58.89 63.89 53.33 61.00 59.28
LongVU 53.41 58.42 50.34 53.20 53.84 55.96 59.52 51.01 58.80 56.32
DivPrune 55.78 61.91 52.28 53.60 55.89 60.85 65.46 52.88 59.40 59.65
Random 55.56 61.41 49.89 53.60 55.12 57.30 63.57 52.13 60.40 58.35
DyCoke 54.37 59.98 51.38 54.60 55.08 59.22 62.60 51.98 63.00 59.20
PruneVid 51.11 58.51 49.66 49.00 52.07 51.41 56.39 49.96 53.00 52.69
STTM 55.26 59.25 50.11 49.40 53.51 57.52 61.73 52.43 60.80 58.12
Scissor 51.89 58.74 51.46 50.00 53.02 56.33 61.68 51.31 62.60 57.98
FastVID 57.19 62.94 52.95 55.00 57.02 - - - - -